Ask your own question, for FREE!
Mathematics 7 Online
OpenStudy (anonymous):

a sharpshooter misses a target 10% of the time. what is the probability that the first miss comes after the 3rd shot? what is the probability that her 2nd miss comes no later than her 10th shot?

OpenStudy (anonymous):

let H and M be hit and miss event respectively then i)P(HHHM)=P(H)*P(H)*P(H)*P(M)=9/10*9/10*9/10*1/10=0.0729 II)P(MHHHHHHHHHM)+P(HMHHHHHHHHM)+P(HHMHHHHHHHM)+....+P(HHHHHHHHHMM) FIND THIS VALUE..AND GETB YOUR ANSWER

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!